OpenStudy (anonymous):

Solve the equation on the interval [0, 2π). tan^2 x sin x = tan^2 x B. pi/2, 2π C. pi/2, π D. 0, π

OpenStudy (anonymous):

@studystagram

OpenStudy (anonymous):

I got D. is that correct

OpenStudy (anonymous):

\[tan^2(x)sin(x)=tan^2(x)\] \[tan^2(x)sin(x)-tan^2(x)\] \[tan^2(x)*(sin(x)-1)=0\]Therefore \[tan^2(x)=0~~or~~sin(x)-1=0\] if\[x=\pi\\tan^2(\pi)=0\] if\[x=\frac{\pi}{2}\\sin\left(\frac{\pi}{2}\right)=1\] but, if\[x=0\\tan^2(0)=0\]The solutions of this equation will be\[S=\left\{0,\frac{\pi}{2},\pi\right\}\]